OpenStudy (hba):

If f(x) = x 2 + 1 and f(g(x)) = 4x – 3, then g(x) =

OpenStudy (hba):

f(x)=x^2+1 and f(g(x)) = 4x – 3, then g(x) =?

OpenStudy (hba):

@hartnn

OpenStudy (hba):

@ash2326

hartnn (hartnn):

what is this ? If f(x) = x 2 + 1

OpenStudy (hba):

i have written the question again ?

OpenStudy (hba):

f(x)=x^2+1 and f(g(x)) = 4x – 3, then g(x) =?

OpenStudy (hba):

@hartnn

hartnn (hartnn):

to get f(g(x)) u replace all "x" in f(x) by g(x) so f(g(x))= g(x)^2+1 = 4x-3 so now find g(x) from g(x)^2+1 = 4x-3

OpenStudy (hba):



hartnn (hartnn):

yes, thats correct.

OpenStudy (hba):

Ok thanks